# VoiceCloak: A Multi-Dimensional Defense Framework against Unauthorized Diffusion-based Voice Cloning

### v2

@@ 4.2. Adversarial Identity Obfuscation
Inspired by prior work showing dedicated speaker embeddings capture identity (Chen et al., 2024), directly manipulating these embeddings offers a direct approach, but its practical utility is hindered by poor transferability when attacking unknown or different encoder models.
@@
Therefore, we explore leveraging speech Representation Learning (Chen et al., 2022b) to extract robust general acoustic representations that inherently capture speaker identity cues yet exhibit reduced downstream dependency. Specifically, we select WavLM (Chen et al., 2022a) as our representation extractor, denoted as $\mathcal{R}(\cdot)$. We then apply adversarial perturbations within this more general feature space to enhance defense transferability. As a baseline approach, we consider maximizing the distance between the perturbed audio $x_{adv}$ and the original one $x_{ref}$ using an untargeted objective:
(5) $\mathcal{L}_{ID}=1-SIM(\mathcal{R}(x_{adv}),\mathcal{R}(x_{ref})),$
@@
where $SIM(x_{1},x_{2})$ represents the cosine similarity metric. Maximizing this effectively pushes $x_{adv}$ away from the reference identity representation domain. However, the untargeted objective lacks specific directionality in high-dimensional space. So we further incorporate a targeted component guided by psychoacoustic principles.
@@
Psychoacoustic studies (Kreiman and Sidtis, 2011; Lavner et al., 2000; Reynolds et al., 2000) indicate that listeners are highly sensitive to specific acoustic cues for speaker identity, with significant perceptual differences often observed between genders (e.g., in F0 and formant structures). Leveraging this, we hypothesize that guiding the adversarial audio’s features towards characteristics of the opposite gender will likely create the strongest perceptual contrast, enhancing identity obfuscation.
@@
(6) $\mathcal{C}_{opp}=\frac{1}{N}\sum_{x_{i}\in\mathcal{X}}\mathcal{R}(x_{i}),$
(7) $\mathcal{L}_{ID}=1-SIM(\mathcal{R}(x_{adv}),\mathcal{R}(x_{ref}))+\underbrace{SIM(\mathcal{R}(x_{adv}),\mathcal{C}_{opp})}_{\text{Gender}},$
where $\mathcal{X}$ is the set of opposite gender utterances and $N$ represents length of $\mathcal{X}$.
@@
Attention Context Divergence.
Building upon the motivation to disrupt conditional guidance via attention for identity obfuscation (Section 3.2), we introduce Attention Context Divergence. This strategy targets the attention mechanism, specifically interfering with how contextual information from the reference audio $x_{ref}$ is utilized.
In DiffVC, conditional information from $x_{ref}$ is integrated via Linear-attention layers (Katharopoulos et al., 2020). Within the U-Net, latent codes representing the content $x_{src}^{t}$ are linearly projected to $Q$ matrix, whereas condition latents $x_{ref}^{t}$ are projected to generate the $K$ and $V$ matrices. Linear-attention first computes a context matrix by aggregating values ($V$) weighted by keys ($K$):
(8) $\mathcal{H}_{ctx}(x_{ref})=Softmax(\phi^{(l,t)}(x_{ref}^{t})W_{K}^{l})(\phi^{(l,t)}(x_{ref}^{t})W_{V}^{l})^{T},$
where $\mathcal{H}_{ctx}$ represents the context hidden state, $\phi^{(l,t)}(\cdot)$ is the deep features of the $l^{th}$ block in U-Net at timestep $t$, and $W_{K}^{l},W_{V}^{l}$ are the projection matrices. Before this context interacts with queries ($Q$) to obtain the attention output: $\mathcal{A}^{l}=(\mathcal{H}_{ctx})^{T}(W_{Q}^{l}\phi^{(l,t)}(x_{src}^{t}))$. This context representation dynamically summarizes the aspects of the reference speaker’s style deemed most relevant for synthesizing the current output part based on content queries. By applying a softmax function, we obtain an explicit probability distribution over sequence positions which represents the distribution of the relevant information. Our strategy, therefore, is to maximize the Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) between the context distribution derived from the original reference and that derived from the adversarial reference:
(9) $\mathcal{L}_{ctx}$ $=D_{KL}(P_{ref}\parallel P_{adv}),$
$\text{where}\ P_{adv}$ $=Softmax(\mathcal{H}_{ctx}(x_{adv})),$
Maximizing this divergence forces the attention focus pattern to deviate from the original, thereby hindering accurate style alignment.
To enhance the impact on identity information, we focus the adversarial perturbation on specific layers within the U-Net’s downsampling path. The rationale is that these earlier layers primarily process coarser structural features and lower-frequency components (Wang et al., 2024) often strongly associated with speaker timbre and identity. Therefore, focusing our proposed loss $\mathcal{L}_{ctx}$ on the attention layers within the U-Net’s downsampling path, the calculation of $P_{adv}$ can be restated as:
(10) $P_{adv}=Softmax(\sum_{l}\mathcal{H}_{ctx}^{l}(x_{adv})),$
@@
4.3. Perceptual Fidelity Degradation
Score Magnitude Amplification.
@@ 4.3. Perceptual Fidelity Degradation
To degrade perceptual fidelity by exploiting the sensitivity of the denoising trajectory, we introduce Score Magnitude Amplification. This design directly interferes with the score function $s_{\theta}$, which is estimated by the U-Net according to Eq. 3 and provides the essential drift term for the reverse SDE. We posit that the magnitude of $s_{\theta}$ relates to the strength of the drift guiding the sample toward the target data manifold. Exploiting this connection, the SMA objective involves maximizing the magnitude of the score prediction $s_{\theta}$:
(11) $\mathcal{L}_{score}=\mathbb{E}_{x_{src}^{t}\sim q(x_{src}^{t}|x_{src}^{0}),t\sim\mathcal{U}(0,T_{adv}]}[\|s_{\theta}(x_{src}^{t},x_{adv}^{t},t)\|_{2}],$
@@
where $q(x_{src}^{t}|x_{src}^{0})$ is forward diffusion distribution, $\mathbb{E}[\cdot]$ calculates the average value, $T_{adv}$ stands for adversarial timesteps which will be discussed in Section 4.4. Optimizing the above formula introduces an erroneous drift strength. Consequently, the denoising trajectory is forcefully steered away from the intended high-quality manifold, resulting in a collapse in the perceptual quality.
@@
Furthermore, the iterative nature of the diffusion process may amplify the impact of this manipulation. Perturbations introduced at earlier timesteps can propagate through subsequent steps. This cumulative effect thus enhances the efficacy of our adversarial strategy.
Noise-Guided Semantic Corruption.
@@
Following the motivation outlined in Section 3.2, we introduce a bidirectional semantic interference strategy. The core idea is twofold: (1) compel the features generated with $x_{adv}$ to diverge from those generated using the original reference $x_{ref}$, and (2) concurrently guide these adversarial features towards a ”semantic-free” target state.
@@
Specifically, consider a layer $l$ within the frozen U-Net $\mathcal{U}_{\theta}$ and timestep $t$, we extract the original features $f^{(l,t)}=\mathcal{U}_{\theta}^{l}(x_{src},x_{ref},t)$ conditioned on $x_{ref}$, and $f^{(l,t)}_{adv}=\mathcal{U}_{\theta}^{l}(x_{src},x_{adv},t)$ corresponding to the adversarial one. Furthermore, to define a ”semantic-free” target, we leverage the U-Net’s activation modes to unstructured information. We extract features $f^{(l,t)}_{noise}=\mathcal{U}_{\theta}^{l}(x_{noise},x_{noise},t)$ by feeding Gaussian white noise $x_{noise}$ as both the source content and the reference condition. $f^{(l,t)}_{noise}$ can be considered to represent unstructured features and lack semantic information. The bidirectional objective aims to maximize the distance between $f^{(l,t)}_{adv}$ and $f^{(l,t)}$ while minimizing it with $f^{(l,t)}_{noise}$. This objective encourages the adversarial features to abandon the original semantic structure and move towards a state of incoherence. The objective is formalized as:
(12) $\mathcal{L}_{sem}=1-\cos(f^{(l,t)}_{adv},f^{(l,t)})+\underbrace{\cos(f^{(l,t)}_{adv},f^{(l,t)}_{noise})}_{\text{Sem-free}},$
@@
4.4. Joint Optimization of Defense Objectives
The final adversarial perturbation $\delta$ for VoiceCloak is optimized to simultaneously achieve identity obfuscation (Section 4.2) and perceptual fidelity degradation (Section 4.3). The joint objectives of this comprehensive defense are formalized as follows
(13) $\mathcal{L}_{total}=$ $(\mathcal{L}_{ID},\mathcal{L}_{ctx},\mathcal{L}_{score},\mathcal{L}_{sem})\Lambda^{T},$
$\delta:=$ $\ \arg\max_{\delta}\ \mathcal{L}_{total},$
@@ 4.4. Joint Optimization of Defense Objectives
The efficacy of our perturbation optimization is influenced by the choice of diffusion timesteps $T_{adv}$ used for gradient computation. Informed by prior work (Yu et al., 2024) indicating early denoising steps primarily reconstruct low-frequency signal structure, we concentrate our optimization on these initial steps to maximize the disruption of fundamental integrity and reduce computational overhead.
Table 1. Comparison of defense effectiveness and adversarial imperceptibility with SOTA methods. Higher values are better for metrics marked with $\uparrow$, and vice versa for those marked with $\downarrow$. The best result is marked in BOLD.
@@
5. Experiments
5.1. Experimental Setup
Datasets. In our experiments, we utilize two primary English speech datasets, including LibriTTS (Zen et al., 2019) and VCTK corpus (Yamagishi et al., 2019). To generate adversarial references ($x_{adv}$), we select subsets of 479 utterances from the LibriTTS dev-clean subset and 500 utterances from VCTK, ensuring approximate gender balance ($50\%$ male/female). Additionally, one male and one female speaker are selected to provide the source content. The evaluation focused on gender-matched conversions.
@@ 5.1. Experimental Setup
Baseline Methods We compare VoiceCloak against existing voice protection methods. Given the lack of protection methods for diffusion-based voice conversion, we adopt the following methods for fair comparison: Attack-VC (Huang et al., 2021), VoicePrivacy (Chen et al., 2024), and VoiceGuard (Li et al., 2023). We also include a naive baseline: adding random Gaussian noise to $x_{ref}$ with an $l_{\infty}$-norm matched to our adversarial perturbations.
Evaluation Metrics We evaluate defense performance using objective metrics and a subjective user study (Sec. 5.2.3). Let $y$ and $y_{adv}$ be the audio synthesized using the original reference $x_{ref}$ and the defended $x_{adv}$, respectively. We use Automatic Speaker Verification (ASV) to quantify identity protection. We measure identity similarity by comparing the embedding of the defended output $y_{adv}$ against the target speaker’s average embedding. The acceptance rate of ASV is then calculated as the proportion of $y_{adv}$ samples whose score $>\tau_{ASV}$. To assess the overall Defense Success Rate (DSR), reflecting both identity obfuscation and quality degradation, we combine the ASV results with a naturalness score predictor NISQA (Mittag and Möller, 2021): successful defense requires failing ASV while also exhibiting low quality ($\mathrm{DS}=s_{ASV}<\tau_{ASV}\land\mathrm{NISQA}(y_{adv})<\tau_{q}$ where we set $\tau_{ASV}=0.25$ and $\tau_{q}=3.0$) Additional objective metrics include Dynamic Time Warping (DTW) (Sakoe and Chiba, 1978) and Structural Similarity Index Measure (SSIM) between $y$ and $y_{adv}$ spectrograms. Perturbation imperceptibility on $x_{adv}$ versus $x_{ref}$ is measured using Perceptual Evaluation of Speech Quality (PESQ) (Rix et al., 2001), Mel-Cepstral Distortion (MCD) (Kubichek, 1993) and Signal-to-Noise Ratio (SNR).
@@
Implementation Details All audio data are uniformly resampled to 22050 Hz. We set the number of optimization iterations to 50 and a step size $\alpha=4\times 10^{-5}$. And $\delta$ is constrained within an $l_{\infty}$-norm ball of $\epsilon=0.002$. We set the adversarial and inference timesteps respectively $T_{adv}=6$ and $T=100$. The loss function combined identity and quality objectives with weights $\Lambda=(1.0,4.5,10,0.85)$. All experiments are conducted on NVIDIA RTX 3090 GPU with a fixed random seed.
5.2. Comparison and Analysis
5.2.1. Comparison with Baselines
@@
Figure 3. Mel spectrograms with $F_{0}$ pitch contours (green lines), and inferred intonation of the corresponding words. Arrows indicate perceived intonation shifts. (Intonation aligns with the ground truth, which is marked by green arrows, and diverges, which is marked by red arrows.)
We conduct a quantitative comparison of our VoiceCloak against baseline methods on the LibriTTS and VCTK datasets, evaluating both defense effectiveness and perturbation imperceptibility. The results are shown in Table 1.
@@
Defense Effectiveness. Our framework demonstrates exceptional efficacy in neutralizing the threat of diffusion-based voice cloning. This is most strikingly evidenced by the Defense Success Rate (DSR)-over 71% on LibriTTS and 63% on VCTK, significantly surpassing all baseline approaches. This success stems from our multi-faceted strategy targeting inherent vulnerabilities within the diffusion process. Specifically, the framework achieves significant identity obfuscation, drastically reducing speaker similarity with ASV down to $11\%$, indicating the failure of voice mimicry. Simultaneously, it ensures substantial perceptual quality degradation, reflected in low NISQA scores, a range typically considered unacceptable generation quality. This dual ability to effectively cripple both identity mimicry and audio usability distinguishes VoiceCloak from defenses that may struggle with diffusion models or focus primarily on one objective. The ability to cripple both identity mimicry and audio usability simultaneously underscores the benefit of our comprehensive design. As expected, the naive Random Noise baseline confirms that unstructured noise provides very limited protection.
@@
Figure 3 further visualizes some disruption results with fundamental frequency ($F_{0}$) and intonations. Applying VoiceCloak protection results in forged speech markedly different from the Undefended forgery. The $F_{0}$ curve is notably degraded, appearing blurred and unpredictable, accompanied by highly inconsistent intonation changes.
Imperceptibility. Regarding perturbation imperceptibility, VoiceCloak achieves performance comparable to most baseline methods, as indicated by similar PESQ, SNR scores, and lowest spectral distortion (MCD). This contrasts with methods like VoiceGuard, which significantly sacrifices audio fidelity without commensurate defense gains. This underscores that directly targeting intrinsic vulnerabilities, rather than simply increasing perturbation magnitude, is a more effective strategy for defending diffusion models while maintaining reasonable imperceptibility.
5.2.2. Protecting Commercial Systems
@@
Figure 4. Protecting commercial speaker verification APIs (Iflytek and Azure) from voice spoofing attacks: average similarity confidence scores (lower scores indicate better defense).
We evaluate our proposed method’s effectiveness in protecting commercial speaker verification (SV) systems, such as Iflytek and Azure platforms, against voice spoofing attacks. These APIs output a confidence score indicating speaker similarity, which higher scores increase the vulnerability to false acceptance from spoofing attacks. Therefore, successful protection aims to minimize this score when verifying protected audio against the original speaker’s enrollment. As shown in Figure 4 our proposed method consistently achieves the lowest confidence scores across both platforms. This indicates that VoiceCloak more effectively decouples the protected audio from the original speaker’s identity, underscoring its superior performance in practical, black-box scenarios.
5.2.3. User Study
@@
Figure 5. User perceptual study results. (a) Timbre Dissimilarity Preference. (b) Corresponding results for perceived Naturalness Disruption.
To assess the human perceptual impact of our approach, we conduct a subjective study with 50 participants evaluating samples from VoiceCloak and baseline methods. Aggregated preferences are shown in Figure 5. The user study results demonstrate the perceptual impact of VoiceCloak. Participants consistently rated audio protected by our method as less similar in timbre to the target speaker and significantly less natural compared to baseline outputs, confirming the framework’s effectiveness in disrupting both identity and quality according to human listeners.
Table 2. Ablation study on the contribution of different settings for Identity Obfuscation. Checkmark ($\checkmark$) indicates the setting is used, while $\downarrow$ signifies that a lower value is more desirable. ”w/o” denotes the exclusion of the specified loss term.
Settings Defense Effectiveness Imperceptibility
@@
- - 1.96 46.82% 0.30 3.66 22.58% 3.34 1.46 34.94
$\checkmark$ - 2.16 8.57% 0.30 3.57 27.74% 3.09 1.45 33.43
w/o Gender - 2.25 19.92% 0.30 3.60 14.40% 3.00 1.36 33.26
- $\checkmark$ 2.31 16.20% 0.27 2.96 62.57% 3.22 1.28 33.58
$\checkmark$ $\checkmark$ 2.13 11.00% 0.27 2.85 69.20% 3.21 1.29 33.48
5.3. Ablation Study
5.3.1. Effectiveness of Adversarial Identity Obfuscation
Table 3. Ablation study on the contribution of different settings for Perceptual Fidelity Degradation. Checkmark ($\checkmark$) indicates the setting is used, while $\downarrow$ signifies that a lower value is more desirable. ”w/o” denotes the exclusion of the specified loss term.
Settings Defense Effectiveness Imperceptibility

### v4

@@ 4.2 Adversarial Identity Obfuscation
Inspired by prior work showing dedicated speaker embeddings capture identity (Chen et al. 2024), directly manipulating these embeddings offers a direct approach, but its practical utility is hindered by poor transferability when attacking unknown or different encoder models.
@@
Therefore, we explore leveraging speech Representation Learning (Chen et al. 2022b) to extract general acoustic representations that inherently capture speaker identity cues. Specifically, we select WavLM (Chen et al. 2022a) as our representation extractor, denoted as $\mathcal{R}(\cdot)$. By applying perturbations within this more general feature space, we aim for broader effectiveness against various models. As a baseline untargeted objective, we consider maximizing the embedding distance:
$\mathcal{L}_{ID}=1-Sim(\mathcal{R}(x_{adv}),\mathcal{R}(x_{ref})),$ (5)
@@
where $Sim(x_{1},x_{2})$ represents the cosine similarity metric. While this untargeted objective effectively pushes $x_{adv}$ away from the reference identity representation domain, it lacks specific directionality in high-dimensional space. So we further incorporate a targeted component guided by psychoacoustic principles.
@@
Psychoacoustic studies (Kreiman and Sidtis 2011; Lavner, Gath, and Rosenhouse 2000) indicate that significant perceptual differences in speaker identity often exist between genders, linked to specific acoustic cues like F0 and formant structures . Leveraging this, we assume that guiding the adversarial embedding towards the opposite gender one will likely create the strongest perceptual contrast, enhancing identity obfuscation.
@@
$\mathcal{C}_{opp}=\frac{1}{N}\sum_{x_{i}\in\mathcal{X}}\mathcal{R}(x_{i}),$ (6)
$\mathcal{L}_{ID}=-Sim(\mathcal{R}_{adv},\mathcal{R}_{ref}+\underbrace{Sim(\mathcal{R}_{adv},\mathcal{C}_{opp})}_{\text{Gender}},$ (7)
where $\mathcal{X}$ is the set of opposite gender utterances, $N$ represents length of $\mathcal{X}$, $\mathcal{R}_{i}$ is the embedding of $x_{i}$ in space $\mathcal{R}$.
@@
Attention Context Divergence
Building upon the motivation to disrupt conditional guidance for identity obfuscation (Section 3.2), we introduce Attention Context Divergence. This strategy targets the attention mechanism to interfere with its use of contextual information from $x_{ref}$.
Commonly in diffusion-based VC, conditional information from $x_{ref}$ is integrated via Linear-attention layers (Katharopoulos et al. 2020). Within the U-Net, latent code representing the content $x_{src}^{t}$ are linearly projected to $Q$ matrix, whereas condition latents $x_{ref}^{t}$ are projected to $K$ and $V$. Linear-attention first computes a context matrix by aggregating values ($V$) weighted by keys ($K$):
$\mathcal{H}_{ctx}(x_{ref})=\text{Softmax}(\phi^{(l,t)}(x_{ref}^{t})W_{K}^{l})(\phi^{(l,t)}(x_{ref}^{t})W_{V}^{l})^{T},$
where $\mathcal{H}_{ctx}$ represents the context hidden state, $\phi^{(l,t)}(\cdot)$ is the deep features of the $l^{th}$ block in U-Net at timestep $t$, and $W_{K}^{l},W_{V}^{l}$ are the projection matrices. Then, the context interacts with queries ($Q$) to obtain the final attention output: $\mathcal{A}^{l}=(\mathcal{H}_{ctx})^{T}(W_{Q}^{l}\phi^{(l,t)}(x_{src}^{t}))$. This context representation provides a dynamic summary of the reference speaker’s stylistic features, weighted by their relevance to the current content queries during synthesis. By applying a softmax function, we obtain an explicit probability distribution over sequence positions which signifies the model’s information focus. Our strategy, therefore, is to maximize the Kullback-Leibler (KL) divergence (Kullback and Leibler 1951) between the context distribution derived from the original reference and the adversarial audio:
$\mathcal{L}_{ctx}$ $=D_{KL}(P_{ref}\parallel P_{adv}),$ (8)
$\text{where}\ P_{adv}$ $=\text{Softmax}(\mathcal{H}_{ctx}(x_{adv})),$
Maximizing this divergence forces the attention pattern to deviate from the original, thereby hindering accurate style alignment.
To enhance the impact on identity, we directs the adversarial pressure on the U-Net’s downsampling path. The rationale is that these earlier layers primarily process coarser, lower-frequency features (Wang et al. 2024) strongly associated with speaker timbre and identity. Therefore, focusing our proposed loss $\mathcal{L}_{ctx}$ on the attention layers within the U-Net’s downsampling path, the calculation of $P_{adv}$ can be restated as:
$P_{adv}=\text{Softmax}(\sum_{l}\mathcal{H}_{ctx}^{l}(x_{adv})),$ (9)
@@
4.3 Perceptual Fidelity Degradation
Score Magnitude Amplification
@@ 4.3 Perceptual Fidelity Degradation
To degrade perceptual fidelity by exploiting the sensitivity of the denoising trajectory, we introduce Score Magnitude Amplification. This design directly interferes with the score function $s_{\theta}$, which is estimated by the U-Net according to Eq. 3 and provides the essential drift term for the reverse SDE. We posit that the magnitude of $s_{\theta}$ relates to the strength of the drift guiding the noisy sample toward the target data manifold. Exploiting this connection, the SMA objective involves maximizing the magnitude of the score prediction $s_{\theta}$:
$\mathcal{L}_{score}=\mathbb{E}_{p_{t}(x),t\sim\mathcal{U}(1,T_{adv})}[\|s_{\theta}(x_{src}^{t},x_{adv}^{t},t)\|_{2}],$ (10)
@@
where $p_{t}(x)$ is the distribution of noisy samples $x^{t}\sim q(x^{t}|x^{0})$, $\mathbb{E}[\cdot]$ calculates the average value, $T_{adv}$ stands for adversarial timesteps which will be discussed in Section 4.4. Optimizing the above formula introduces an erroneous drift strength. Consequently, the denoising trajectory is forcefully diverted, resulting in a collapse in perceptual quality.
@@
Furthermore, the iterative nature of the diffusion process may amplify these induced trajectory deviations. Perturbations introduced at earlier timesteps can propagate through subsequent steps. This error cumulative effect thus enhances the efficacy of our adversarial strategy.
Noise-Guided Semantic Corruption
@@
Following the motivation outlined in Section 3.2, we introduce a bidirectional semantic interference strategy. The core idea is twofold: (1) compel the features generated with $x_{adv}$ to diverge from those generated using the original reference $x_{ref}$, and (2) concurrently guide these adversarial features towards a ”semantic-free” state.
@@
Specifically, consider a network layer $l$ within the frozen U-Net $\mathcal{U}_{\theta}$ and timestep $t$, we extract the original features $f^{(l,t)}=\mathcal{U}_{\theta}^{l}(x_{src},x_{ref},t)$ conditioned on $x_{ref}$, and $f^{(l,t)}_{adv}=\mathcal{U}_{\theta}^{l}(x_{src},x_{adv},t)$ corresponding to the adversarial version. Furthermore, to define a ”semantic-free” target, we leverage the U-Net’s activation modes to unstructured information. We extract features $f^{(l,t)}_{noise}=\mathcal{U}_{\theta}^{l}(x_{noise},x_{noise},t)$ by feeding Gaussian white noise $x_{noise}$ as both the source content and the reference condition. $f^{(l,t)}_{noise}$ can be considered to represent unstructured features and lack semantic information. The bidirectional objective aims to maximize the distance between $f^{(l,t)}_{adv}$ and $f^{(l,t)}$ while minimizing it with $f^{(l,t)}_{noise}$. This objective encourages the adversarial features to abandon the original semantic structure and move towards a state of incoherence which can be formalized as:
$\mathcal{L}_{sem}=1-\cos(f^{(l,t)}_{adv},f^{(l,t)})+\underbrace{\cos(f^{(l,t)}_{adv},f^{(l,t)}_{noise})}_{\text{Sem-free}},$ (11)
@@
4.4 Joint Optimization of Defense Objectives
The final adversarial perturbation $\delta$ for VoiceCloak is optimized to simultaneously achieve our dual objectives (Section 4.2 and 4.3). The joint objectives of this comprehensive defense are formalized as follows
$\mathcal{L}_{total}=$ $(\mathcal{L}_{ID},\mathcal{L}_{ctx},\mathcal{L}_{score},\mathcal{L}_{sem})\Lambda^{T},$ (12)
$\delta$= $\ \arg\max_{\delta}\ \mathcal{L}_{total},$
@@ 4.4 Joint Optimization of Defense Objectives
The efficacy of our perturbation optimization is influenced by the choice of diffusion timesteps $T_{adv}$ used for gradient computation. Informed by prior work (Yu et al. 2024) indicating early denoising steps primarily reconstruct low-frequency overall structural signal, we concentrate the optimization on these initial steps to maximize the disruption of fundamental integrity and reduce computational overhead.
@@
Table 1: Comparison of defense effectiveness and adversarial imperceptibility with SOTA methods. Higher values are better for metrics marked with $\uparrow$, and vice versa for those marked with $\downarrow$. The best result is marked in BOLD.
5 Experiments
5.1 Experimental Setup
Datasets
Experiments are conducted on the LibriTTS (Zen et al. 2019) and VCTK (Yamagishi, Veaux, and MacDonald 2019) datasets. We selected a gender-balanced audio subsets (479 utterances from LibriTTS, 500 from VCTK) to generate adversarial reference speech set $\mathcal{D}_{x_{adv}}$.
Baseline Methods
@@ 5.1 Experimental Setup
We compare VoiceCloak against existing voice protection methods. We adopt the following methods for fair comparison: Attack-VC (Huang et al. 2021), VoicePrivacy (Chen et al. 2024), and VoiceGuard (Li et al. 2023). We also include a naive baseline: adding random Gaussian noise to $x_{ref}$.
Evaluation Metrics
For identity protection, we report the Automatic Speaker Verification (ASV) acceptance rate, where a lower rate for protected outputs signifies more effective obfuscation. We define a comprehensive Defense Success Rate (DSR) to measure the achievement of both our objectives. A defense is considered successful if the protected output both fails speaker verification and exhibits low perceptual quality ($\mathrm{DS}=s_{ASV}<\tau_{ASV}\land\mathrm{NISQA}(y_{adv})<\tau_{q}$), with thresholds $\tau_{ASV}=0.25$ and $\tau_{q}=3.0$. Additional metrics include Dynamic Time Warping (DTW) (Sakoe and Chiba 1978) and SSIM between $y$ and $y_{adv}$ spectrograms. Perturbation imperceptibility on $x_{adv}$ versus $x_{ref}$ is measured using PESQ (Rix et al. 2001), Mel-Cepstral Distortion (MCD) (Kubichek 1993) and SNR.
Implementation Details
@@
We conduct our experiments mainly using DiffVC (Popov et al. 2022) as the target system. We set the number of optimization iterations to 50 and a step size $\alpha=4\times 10^{-5}$. And $\delta$ is constrained within an $l_{\infty}$-norm ball of $\epsilon=0.002$. We set the adversarial and inference timesteps respectively $T_{adv}=6$ and $T=100$. The loss function combined identity and quality objectives with weights $\Lambda=(1.0,4.5,10,0.85)$. All experiments are conducted on NVIDIA RTX 3090 GPU with a fixed random seed.
5.2 Comparison and Analysis
Comparison with Baselines
@@ 5.2 Comparison and Analysis
Figure 3: Mel spectrograms with $F_{0}$ pitch contours (green lines), and inferred intonation of the corresponding words. Arrows indicate perceived intonation shifts. (Intonation aligns with the ground truth, which is marked by green arrows, and diverges, which is marked by red arrows.)
@@
As shown in Table 1, VoiceCloak demonstrates exceptional defense efficacy, significantly outperforming all baselines with a Defense Success Rate (DSR) of over 71% on LibriTTS and 63% on VCTK. This high DSR reflects success in both of our defense objectives: identity obfuscation is achieved by drastically reducing speaker verification acceptance rates to  11%, while perceptual quality degradation is confirmed by low NISQA scores indicating unacceptable audio generation quality. This dual ability to effectively cripple both identity mimicry and audio usability distinguishes VoiceCloak from defenses that may struggle with diffusion models or focus primarily on one objective. As expected, the naive Random Noise baseline confirms that unstructured noise provides very limited protection. Regarding imperceptibility, VoiceCloak performs comparably to baselines. This supports our strategy of targeting intrinsic vulnerabilities, rather than simply increasing perturbation magnitude.
@@
Figure 3 further visualizes disruption result. Applying VoiceCloak protection results markedly different from the undefended one. The $F_{0}$ curve is notably degraded, appearing blurred and unpredictable, accompanied by highly inconsistent intonation changes.
Protecting Commercial Systems
@@
Figure 4: Protecting commercial speaker verification APIs (Iflytek, Azure) from spoofing attacks (lower are better).
We also evaluated the effectiveness in protecting commercial speaker verification (SV) APIs (Iflytek, Azure) to simulate real-world anti-spoofing scenarios. Successful protection aims to minimize the similarity score returned by the API. Figure 4 demonstrated VoiceCloak’s superior ability to decouple the protected audio from the original speaker’s identity.
User Study
@@
Figure 5: User perceptual study results. (a) Timbre Dissimilarity Preference. (b) Corresponding results for perceived Naturalness Disruption.
We conducted a user study with 50 participants to assess perceptual impact. Listeners performed comparisons on two criteria: Timbre Dissimilarity and Naturalness Disruption and we aggregated results in Figure 5, where ”Neutral” indicates no perceived difference. Participants consistently rated VoiceCloak’s outputs as having both greater timbre dissimilarity and more severe naturalness disruption, confirming its human-perceived effectiveness.
Settings Defense Effectiveness
@@
- - 1.96 46.82% 0.30 3.66 22.58%
$\checkmark$ - 2.16 8.57% 0.30 3.57 27.74%
w/o Gender - 2.25 19.92% 0.30 3.60 14.40%
- $\checkmark$ 2.31 16.20% 0.27 2.96 62.57%
$\checkmark$ $\checkmark$ 2.13 11.00% 0.27 2.85 69.20%
Table 2: Ablation study on the contribution of different settings for Identity Obfuscation. $\checkmark$ indicates the setting is used, ”w/o” denotes the exclusion of the specified term.
5.3 Ablation Study
Effectiveness of Adversarial Identity Obfuscation
Settings Defense Effectiveness